OpenStudy (anonymous):

solve give exact solution (r+5)^2=10

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow r + 5 = \sqrt{10} }\) \(\Large \color{MidnightBlue}{\Rightarrow r = \sqrt{10} -5 }\)

OpenStudy (anonymous):

thanks

Parth (parthkohli):

:) can you attempt the other questions on your own?

Parth (parthkohli):

Because I'd really love you to learn, rather than just ask.

OpenStudy (anonymous):

i am trying but I am having the hardest time getting this